OpenStudy (anonymous):

identify the conic (y-4)^2=16(x+1)

OpenStudy (sidsiddhartha):

parabola is'nt it

OpenStudy (anonymous):

It is a parabola

OpenStudy (anonymous):

We have 2 different types of equation for a parabola \[(x-k)^2=2p*(y-h)\] or \[(y-k)^2=2p*(x-h)\] This parabola is the second type, so \[(y-4)^2=16*(x+1)\] Is a Parabola.